### v3

# Pivot-minors and the Erdős-Hajnal conjecture

Abstract
We prove a conjecture of Kim and Oum that every proper pivot-minor-closed class of graphs has the strong Erdős-Hajnal property. More precisely, for every graph $H$, there exists $\epsilon>0$ such that every $n$-vertex graph with no pivot-minor isomorphic to $H$ contains two sets $A,B$ of vertices such that $|A|,|B|\geqslant\epsilon n$ and $A$ is complete or anticomplete to $B$.
1 Introduction
For a graph $G$, we let $\alpha(G),\omega(G)$ denote the independence and clique number of $G$ respectively. A class of graph $\mathcal{G}$ has the Erdős-Hajnal property if there exists a constant $c>0$ such that $\max\{\alpha(G),\omega(G)\}\geqslant|V(G)|^{c}$ for every $G\in\mathcal{G}$. Erdős and Hajnal [5] conjectured that for every graph $H$, the graphs with no induced $H$ subgraph have the Erdős-Hajnal property.
Strengthening the Erdős-Hajnal property is the strong Erdős-Hajnal property. Two vertex sets $A,B\subseteq V(G)$ in a graph $G$ are complete if $A,B$ are disjoint, and every vertex of $A$ is adjacent to every vertex of $B$. Similarly, $A$ and $B$ are anticomplete if they are disjoint and no vertex of $A$ is adjacent to a vertex of $B$. A class of graphs $\mathcal{G}$ has the strong Erdős-Hajnal property if every graph $G\in\mathcal{G}$ with $|V(G)|>1$ contains two sets $A,B$ of vertices such that $|A|,|B|\geqslant\epsilon|V(G)|$ and $A$ is either complete or anticomplete to $B$. The strong Erdős-Hajnal property implies the Erdős-Hajnal property, however many classes, such as triangle-free graphs do not satisfy the strong Erdős-Hajnal property [7].
Chudnovsky and Oum [1] proved an analogue of the Erdős-Hajnal conjecture that for every graph $H$, the class of $H$-vertex-minor-free graphs has the (strong) Erdős-Hajnal property. Kim and Oum [8] conjectured a strengthening that for every graph $H$, the class of $H$-pivot-minor-free graphs has the (strong) Erdős-Hajnal property. We prove this conjecture. Pivot-minors shall be formally defined in Section 2.
Theorem 1.1.
For every graph $H$, there exists $\epsilon>0$ such that for all $n>1$, every $n$-vertex graph with no pivot-minor isomorphic to $H$ has two sets $A,B$ of vertices such that $|A|,|B|\geqslant\epsilon n$ and $A$ is either complete or anticomplete to $B$.
Since the strong Erdős-Hajnal property implies the Erdős-Hajnal property, we obtain the following corollary of Theorem 1.1.
Corollary 1.2.
For every graph $H$, there exists $c>0$ such that $\max\{\alpha(G),\omega(G)\}\geqslant|V(G)|^{c}$ for every $H$-pivot-minor-free graph $G$.
Theorem 1.1 was proved in the case that $H$ is a cycle by Kim and Oum [8]. Two other natural pivot-minor-closed classes are circle graphs and graphs of bounded rank-width. Both of these classes are also vertex-minor-closed, and so have the strong Erdős-Hajnal property by the theorem of Chudnovsky and Oum [1]. The strong Erdős-Hajnal property for circle graphs is also implied by an earlier theorem of Pach and Solymosi [13]. Almost tight bounds for the Erdős-Hajnal property of circle graphs are now known. A theorem of the author [4] implies that $\max\{\alpha(G),\omega(G)\}\geqslant\frac{1}{3}\sqrt{|V(G)|/\log_{2}|V(G)|}$ for every circle graph $G$. Up to a constant factor, this is best possible due to a construction of Kostochka [9, 10].
Recently the strong Erdős-Hajnal property has been studied for other hereditary classes of graphs. For a graph $H$, we let $\overline{H}$ denote the complement of $H$. Chudnovsky, Scott, Seymour, and Spirkl [2] proved that for every pair of trees $T,F$, the class of graphs containing no induced $T$ or $\overline{F}$ subgraph has the strong Erdős-Hajnal property.
A subdivision of a graph $G$ is a graph obtainable from $G$ by replacing edges by possibly longer paths. Chudnovsky, Scott, Seymour, and Spirkl [3] also proved that for every graph $H$, the class of graphs containing no subdivision of $H$, or the complement of a subdivision of $H$ as an induced subgraph has the strong Erdős-Hajnal property. We shall consider a more general class of graphs.
A proper odd subdivision of a graph $G$ is a graph obtainable from $G$ by replacing each edge by a path of odd length at least $3$. A fuzzy odd path is a graph obtained from an odd length path by possibly adding an additional edge $uv$ between two internal vertices of the path, such that $uv$ is contained in a triangle. We say that a graph $G$ is a proper fuzzy odd subdivision of a graph $H$ if $G$ can be obtained from $H$ by replacing each edge with a fuzzy odd path of length more than one. To find a $r$-vertex graph $H$ as a pivot-minor, it is enough to instead find as an induced subgraph either a proper fuzzy odd subdivision of $K_{r}$ (see Lemma 2.2) or the complement of a graph obtainable from $K_{r+1}$ by replacing each edge with a longer path of length $1\pmod{3}$ (see Lemma 2.5). By finding induced proper fuzzy odd subdivisions, we actually also prove the following theorem, which generalizes the theorem of Chudnovsky, Scott, Seymour and, Spirkl [3] on induced subdivisions and their complements.
Theorem 1.3.
For every graph $H$, there exists $\epsilon>0$ such that for all $n>1$, every $n$-vertex graph containing no proper fuzzy odd subdivision of $H$ or the complement of a fuzzy odd subdivision of $H$ as an induced subgraph, contains two sets $A,B$ of vertices such that $|A|,|B|\geqslant\epsilon n$ and $A$ is either complete or anticomplete to $B$.
Our proof of Theorem 1.1 and Theorem 1.3 adapts and extends that of Chudnovsky, Scott, Seymour, and Spirkl [3]. It would be natural to conjecture a further strengthening of Theorem 1.3. Namely, that Theorem 1.3 holds simply for odd subdivisions of $H$, rather than fuzzy odd subdivisions. However, this turns out to be false. For sufficiently large $n$, Fox [6] constructed $n$-vertex graphs containing no odd hole or antihole and with no complete or anticomplete pair $A,B\subseteq V(G)$ with $|A|,|B|\geqslant\frac{15n}{\log_{2}n}$. However, Scott, Seymour, and Spirkl [15] did prove that for every graph $H$ and $c>0$, there exists $\epsilon>0$ such that for all $n>1$, every $n$-vertex graph containing no odd subdivision of $H$ or the complement of an odd subdivision of $H$ as an induced subgraph, has two sets $A,B$ of vertices such that $|A|\geqslant\epsilon n$, $|B|\geqslant\epsilon n^{1-c}$, and $A$ is either complete or anticomplete to $B$. If instead of forbidding odd subdivision, one instead forbids induced subdivisions where each edge is replaced with a longer path with length $2\ell\pmod{k}$, then these graphs do have the strong Erdős-Hajnal property (see Theorem 3.6).
In Section 2 we formally introduce pivot-minor and prove all the necessary lemmas concerning them. The rest of the paper will then focus on induced subgraphs. In Section 3 there is a further overview of the remainder of the proofs of Theorem 1.1 and Theorem 1.3.
2 Pivot-minors and subdivisions
For a vertex $v$ in a graph $G$, we let $N(v)$ denote its neighbourhood, i.e., the set of vertices adjacent to $v$. We let $N[v]=N(v)\cup\{v\}$. When the graph $G$ is not clear from context, we use $N_{G}(v),N_{G}[v]$.
For an edge $uv$ of a graph $G$, let $V_{1}=N(u)\backslash N[v]$, $V_{2}=N(v)\backslash N[u]$, and $V_{3}=N(u)\cap N(v)$. Pivoting the edge $uv$ of $G$ is the act of first complementing the edges between each of the three pairs of vertex sets $(V_{1},V_{2})$, $(V_{2},V_{3})$, and $(V_{1},V_{3})$, and then swapping the vertex labels of $u$ and $v$. We denote this graph by $G\wedge uv$. A graph $H$ is a pivot-minor of a graph $G$ if $H$ can be obtained from $G$ by a sequence of vertex deletions and pivots.
Given a graph $H$ and an integer $t\geqslant 1$, we denote by $H^{t}$ the subdivision of $H$ obtained by replacing each edge with a path of length $t+1$. For a positive integer $r$, we let $[r]=\{1,\ldots,r\}$. For a graph $G$ and $X\subseteq V(G)$, we let $G[X]$ denote the induced subgraph of $G$ on vertex set $X$.
We begin with a simple universal graph for pivot-minors.
Lemma 2.1.
The graph $K_{r}^{2}$ contains every graph on at most $r$ vertices as a pivot-minor.
Proof.
Clearly it is enough to prove that $K_{r}^{2}$ contains every graph on exactly $r$ vertices as a pivot-minor. So, let $H$ be a graph on vertex set $[r]$. Let $V(K_{r}^{2})=\{x_{1},\ldots,x_{r}\}\cup\bigcup_{1\leqslant i<j\leqslant r}\{y_{i,j},z_{i,j}\}$, where for every pair $1\leqslant i<j\leqslant r$, $x_{i}y_{i,j}z_{i,j}x_{j}$ is a path. Now, we find a pivot-minor isomorphic to $H$ by pivoting the edge $y_{i,j}z_{i,j}$ for every $ij\in E(H)$ with $i<j$, and then deleting the vertex set $\bigcup_{1\leqslant i<j\leqslant r}\{y_{i,j},z_{i,j}\}$. ∎
This can be extended to proper fuzzy odd subdivisions.
Lemma 2.2.
Every proper fuzzy odd subdivision of $K_{r}$ contains every graph on at most $r$ vertices as a pivot-minor.
Proof.
Let $H$ be a proper odd subdivision of $K_{r}$. If $H=K_{r}^{2}$, then the lemma follows by Lemma 2.1. So we argue inductively on $|V(H)|$. If $|V(H)|>|V(K_{r}^{2})|$, then $H$ contains vertices $w,x,y,z$ such that either $wxyz$ is an induced path and $w,x,y,z$ all have degree 2 in $H$, or $N(x)=\{w,y,z\}$, $N(y)=\{x,z\}$, $w$ has degree 2 in $H$, and $z$ has degree 3 in $H$. In either case, $H^{\prime}=(H\wedge xy)-\{x,y\}$ is a proper fuzzy odd subdivision of $K_{r}^{2}$ with $|V(K_{r}^{2})|\leqslant|V(H^{\prime})|=|V(H)|-2$. So, $H^{\prime}$, (and thus $H$) contains every graph on at most $r$ vertices as a pivot-minor. ∎
For two graphs $G,H$, we let $G+H$ denote the graph obtained from the disjoint union of $G$ and $H$ by adding an edge between every vertex of $G$ and every vertex of $H$.
Lemma 2.3.
The graph $\overline{K_{r}^{3}}+K_{1}$ contains every graph on at most $r$ vertices as a pivot-minor.
Proof.
Clearly it is enough to prove that $\overline{K_{r}^{3}}+K_{1}$ contains every graph on exactly $r$ vertices as a pivot-minor. So, let $H$ be a graph on vertex set $[r]$. Let $V(\overline{K_{r}^{3}}+K_{1})=\{x_{1},\ldots,x_{r}\}\cup\bigcup_{1\leqslant i<j\leqslant r}\{x_{i,j}^{1},x_{i,j}^{2},x_{i,j}^{3}\}\cup\{u\}$, where for every pair $1\leqslant i<j\leqslant r$, $x_{i}x_{i,j}^{1}x_{i,j}^{2}x_{i,j}^{3}x_{j}$ is a path in $K_{r}^{3}$, and $u$ is the vertex of the $K_{1}$. Now, we find a pivot-minor isomorphic to $H$ by pivoting the edge $ux_{i,j}^{2}$ and then the edge $x_{i,j}^{1}x_{i,j}^{3}$ for every $ij\notin E(H)$ with $i<j$, and then deleting the vertex set $\bigcup_{1\leqslant i<j\leqslant r}\{x_{i,j}^{1},x_{i,j}^{2},x_{i,j}^{3}\}\cup\{u\}$. ∎
Lemma 2.4.
Let $H$ be a graph obtained from $K_{r}$ by replacing each edge with a longer path of length $1\pmod{3}$. Then, the graph $\overline{H}+K_{1}$ contains every graph on at most $r$ vertices as a pivot-minor.
Proof.
If $H=K_{r}^{3}$, then the result follows from Lemma 2.3. So, we shall proceed inductively on $|V(H)|$. We may assume that $|V(H)|>|V(K_{r}^{3})|$. Therefore, there exists an induced path $x_{1}x_{2}x_{3}x_{4}x_{5}$ of $H$ such that $x_{2},x_{3},x_{4}$ each have degree 2 in $H$ and $H^{\prime}=(H\backslash\{x_{2},x_{3},x_{4}\})\cup\{x_{1}x_{5}\}$ is a graph obtained from $K_{r}$ by replacing each edge with a longer path of length $1\pmod{3}$. Since $|V(H^{\prime})|<|V(H)|$, by inductive hypothesis, it follows that $\overline{H^{\prime}}+K_{1}$ contains every graph on at most $r$ vertices as a pivot-minor. Therefore, it is enough to show that $\overline{H}+K_{1}$ contains $\overline{H^{\prime}}+K_{1}$ as a pivot-minor. Let $u$ be the vertex of the $K_{1}$ in $\overline{H}+K_{1}$. Then observe that $((\overline{H}+K_{1})\wedge ux_{3}\wedge x_{2}x_{4})\backslash\{x_{2},x_{3},x_{4}\}=\overline{H^{\prime}}+K_{1}$, as desired. ∎
Immediately from this lemma, we obtain the following slightly more convenient lemma.
Lemma 2.5.
Let $H$ be a graph obtained from $K_{r+1}$ by replacing each edge with a longer path of length $1\pmod{3}$. Then, the graph $\overline{H}$ contains every graph on at most $r$ vertices as a pivot-minor.
Let us remark that Lemma 2.5 actually highlights one difference in the structure of pivot-minor-closed classes of graphs compared to vertex-minor-closed classes of graphs. The complement $\overline{G}$ of a graph $G$ is a special case of a rank-$1$ perturbation of $G$. Vertex-minors are well behaved under bounded rank perturbations. To be more precise, if $\mathcal{F}$ is a proper vertex-minor-closed class of graphs, then not every graph can be obtained as a vertex-minor of a bounded rank perturbation of a graph in $\mathcal{F}$ (see [11, Lemma 1.6.7]). However, this turns out to not be the case for pivot-minors.
Bipartite graphs are a proper pivot-minor-closed class of graphs and for each $r$, $K_{r+1}^{3}$ is bipartite, in particular $\overline{K_{r+1}^{3}}$ is a rank-1 perturbation of a bipartite graph. Therefore, by Lemma 2.5, there exists proper pivot-minor-closed classes of graphs whose rank-1 perturbations contain all graphs as pivot-minors. Another natural proper pivot-minor-closed class of graphs are the pivot-minors of line graphs [12]. As pointed out by Sang-il Oum, this class also contains $K_{r+1}^{3}$, thus this class also has the property that its rank-1 perturbations contain all graphs as pivot-minors.
We require another lemma on induced subgraphs that shall allow to more easily apply Lemma 2.5.
Let $G$ be a graph and let $F$ be a subgraph of $G$. Let $H$ be a graph obtained from $G$ by subdividing at least once every edge of $G$ not in $E(F)$, and not subdividing the edges in $E(F)$. We call such a graph $H$ (and graphs isomorphic to it) a $F$-filleting of $G$. Chudnovsky, Scott, Seymour, and Spirkl [3] observed that for every pair of positive integers $\ell,k$, there exists a graph $G$ and a subgraph $F$ that is a forest whose components are all paths, such that every $F$-filleting of $G$ contains a hole of length $2\ell\pmod{k}$. We shall require a slight extension for induced subdivisions rather than holes.
Lemma 2.6.
Let $\ell,k,r$ be a triple of positive integers. Then there exists a graph $G$ with a subgraph $F$ that is a forest whose components are all paths, such that every $F$-filleting of $G$ contains as an induced subgraph a graph $J$ that is obtainable from $K_{r}$ by replacing each edge with a longer path of length $2\ell\pmod{k}$.
Proof.
Let $R=R_{k^{2}}(r+3kr^{2})$ be the multicolour Ramsey number such that every $k^{2}$-edge-colouring of $K_{R}$ contains a monochromatic $K_{r+3kr^{2}}$. Now let $G$ be obtained from two disjoint copies of $K_{R}$ on vertex sets $\{x_{1},\ldots,x_{R}\}$ and $\{y_{1},\ldots,y_{R}\}$ by for each $1\leqslant i\leqslant R$ adding a path $F_{i}$ between $x_{i}$ and $y_{i}$ of length $\ell$. Let $F$ be the union of these paths.
Let $H$ be some $F$-filleting of $G$. For each pair $1\leqslant i<j\leqslant R$, let $X_{i,j}$ be the path in $H$ corresponding to the edge $x_{i}x_{j}$ of $G$, and similarly, let $Y_{i,j}$ be the path in $H$ corresponding to the edge $y_{i}y_{j}$ of $G$. By Ramsey’s theorem, there exists some $T\subset R$ with $|T|=r+3kr^{2}$ and some pair $0\leqslant\ell_{x},\ell_{y}<k$ such that for every pair $i,j\in T$ with $i<j$, we have that $X_{i,j}$ is a path of length $\ell_{x}\equiv\pmod{k}$ and $Y_{i,j}$ is a path of length $\ell_{y}\equiv\pmod{k}$. Without loss of generality, we may assume that $T=[r+3kr^{2}]$.
Now for each pair $1\leqslant i<j\leqslant r$, let $P_{i,j}^{1}=X_{i,r+3k(ir+j)+1}\cup\bigcup_{t=1}^{k-1}X_{r+3k(ir+j)+t,r+3k(ir+j)+t+1}$, $P_{i,j}^{2}=\bigcup_{t=0}^{k-1}Y_{r+3k(ir+j)+k+t,r+3k(ir+j)+k+t+1}$, $P_{i,j}^{3}=\bigcup_{t=0}^{k-2}X_{r+3k(ir+j)+2k+t,r+3k(ir+j)+2k+t+1}\cup X_{r+3k(ir+j)+3k-1,j}$, and let $P_{i,j}=P_{i,j}^{1}\cup F_{r+3k(ir+j)+k}\cup P_{i,j}^{2}\cup F_{r+3k(ir+j)+2k}\cup P_{i,j}^{3}$. Notice that each path $P_{i,j}$ has length $2\ell\pmod{k}$. Then, taking $J$ to be the union of all such paths $P_{i,j}$ provides the desired induced subgraph of $G$. ∎
By Lemma 2.5 and Lemma 2.6, we obtain the following.
Lemma 2.7.
For every graph $H$, there exists a graph $G$ with a subgraph $F$ that is a forest whose components are all paths, such that the complement of every $F$-filleting of $G$ contains $H$ as a pivot-minor.
3 Coherence
In this section we shall introduce some required preliminaries and show that Theorem 1.1 and Theorem 1.3 are implied by Theorem 3.1, which in turn is implied by two lemmas, which each of the following two sections are then dedicated to proving. As in [3], our results apply more generally to massed graphs, which we introduce next.
A function $\mu:2^{V(G)}\to\mathbb{R}$ is a mass on a graph $G$ if
  • $\mu(\emptyset)=0$ and $\mu(V(G))=1$,
  • $\mu(X)\leqslant\mu(Y)$ for all $X\subseteq Y\subseteq V(G)$, and
  • $\mu(X\cup Y)\leqslant\mu(X)+\mu(Y)$ for all $X,Y\subseteq V(G)$.
We call such a pair $(G,\mu)$ a massed graph. The typical example of a mass on a graph $G$ is $\mu(X)=|X|/|V(G)|$. Another example is $\mu(X)=\chi(G[X])$.
For $\epsilon>0$, let us say that a massed graph $(G,\mu)$ is $\epsilon$-coherent if
  • $\mu(v)<\epsilon$ for every $v\in V(G)$,
  • $\mu(N(v))<\epsilon$ for every $v\in V(G)$, and
  • $\min\{\mu(A),\mu(B)\}<\epsilon$ for every anticomplete pair $A,B\subseteq V(G)$.
We will prove the following.
Theorem 3.1.
For every graph $H$, there exists $\epsilon>0$ such that every $\epsilon$-coherent massed graph $(G,\mu)$ contains a proper fuzzy odd subdivision of $H$ as an induced subgraph.
By Lemma 2.2, we also obtain the following corollary of Theorem 3.1.
Corollary 3.2.
For every graph $H$, there exists $\epsilon>0$ such that every $\epsilon$-coherent massed graph $(G,\mu)$ contains a pivot-minor isomorphic to $H$.
The following is Rödl’s [14] theorem.
Theorem 3.3 (Rödl [14]).
For every graph $H$ and $\epsilon>0$, there exists $\delta>0$ such that for every graph $G$ containing no induced $H$ subgraph, there exists $X\subseteq V(G)$ with $|X|\geqslant\delta|V(G)|$ such that in one of $G[X]$, $\overline{G[X]}$, every vertex in $X$ has degree less than $\epsilon|X|$.
With Rödl’s [14] theorem, we can prove Theorem 1.3 assuming Theorem 3.1. This is just the standard argument of applying Rödl’s [14] theorem, which is often used to prove the strong Erdős-Hajnal property by splitting into a sparse case and a dense case.
Proof of Theorem 1.3 assuming Theorem 3.1..
For our graph $H$, choose $\epsilon^{\prime}>0$ satisfying Theorem 3.1. Let $H^{\prime}$ be some proper fuzzy odd subdivision of $H$. Now choose $\delta>0$ satisfying Theorem 3.3 for $H^{\prime},\epsilon^{\prime}$. Now, let $\epsilon=\epsilon^{\prime}\delta$.
Let $G$ be a graph containing no proper fuzzy odd subdivision of $H$ or the complement of a proper fuzzy odd subdivision of $H$ as an induced subgraph, and with $|V(G)|>1$. It now follows from Theorem 3.3 that there exists $X\subseteq V(G)$ with $|X|\geqslant\delta|V(G)|$ such that in one of $G[X]$, $\overline{G[X]}$, every vertex in $X$ has degree less than $\epsilon^{\prime}|X|$. If every in $X$ has degree less than $\epsilon^{\prime}|X|$ in $G[X]$, then let $G^{*}=G[X]$, otherwise let $G^{*}=\overline{G[X]}$. For each $Y\subseteq X$, let $\mu(Y)=\frac{|Y|}{|X|}$. Then $(G^{*},\mu)$ is a massed graph. Since $G$ contains no proper fuzzy odd subdivision of $H$ or the complement of a proper fuzzy odd subdivision of $H$ as an induced subgraph, we may assume that $G^{*}$ contains no proper fuzzy odd subdivision of $H$ as an induced subgraph. So by Theorem 3.1, $G^{*}$ is not $\epsilon^{\prime}$-coherent.
By definition, we have that $\mu(N(v))=|N(v)|/|X|<\epsilon^{\prime}|X|/|X|=\epsilon^{\prime}$ for each $v\in X$. If $\frac{1}{|X|}\geqslant\epsilon^{\prime}$, then $|V(G)|\leqslant\frac{1}{\epsilon^{\prime}\delta}=\frac{1}{\epsilon}$, and so there exists distinct vertices $u,v\in V(G)$ with $A=\{u\},B=\{v\}$ either complete or anticomplete, and $|A|=|B|=1\geqslant\epsilon|V(G)|$. So, we may assume that $\frac{1}{|X|}<\epsilon^{\prime}$. Therefore, $\mu(v)=1/|X|<\epsilon^{\prime}$ for all $v\in X$. Since $G^{*}$ is not $\epsilon^{\prime}$-coherent, it follows that there exists an anticomplete pair $A,B\subseteq X$ with $\mu(A),\mu(B)\geqslant\epsilon^{\prime}$. Therefore, $A,B$ are either complete or anticomplete in $G$, and $|A|,|B|\geqslant\epsilon^{\prime}|X|\geqslant\epsilon^{\prime}\delta|V(G)|\geqslant\epsilon|V(G)|$, as desired. ∎
As seen from the above proof, Rödl’s [14] theorem splits proving the strong Erdős-Hajnal property into handling a sparse and a dense case. Corollary 3.2 handles the sparse case for pivot-minors. For the dense case we require the following theorem of Chudnovsky, Scott, Seymour, and Spirkl [3], which is a technical strengthening of their result that for every graph $H$, the class of graphs containing no subdivision of $H$, or the complement of a subdivision of $H$ as an induced subgraph has the strong Erdős-Hajnal property.
Theorem 3.4 (Chudnovsky, Scott, Seymour, and Spirkl [3, 2.3]).
For every graph $H$ and subgraph $F$ that is a forest whose components are all paths, there exists $\epsilon>0$ such that every $\epsilon$-coherent massed graph $(G,\mu)$ contains a $F$-filleting of $H$ as an induced subgraph.
By Lemma 2.7, Theorem 3.4 then handles the dense case for pivot-minors. Thus, to prove Theorem 1.1, it now just remains to prove Theorem 3.1 (as this implies Corollary 3.2 which handles the sparse case).
Let us remark that by Lemma 2.6, one can also obtain the following two theorems from Theorem 3.4 and then from Rödl’s [14] theorem (in the same way as in the proof of Theorem 1.3 assuming Theorem 3.1).
Theorem 3.5.
For every graph $H$, there exists $\epsilon>0$ such that every $\epsilon$-coherent massed graph $(G,\mu)$ contains as an induced subgraph a graph $J$ that is obtainable from $H$ by replacing each edge with a longer path of length $2\ell\pmod{k}$.
Theorem 3.6.
For every graph $H$, there exists $\epsilon>0$ such that for all $n>1$, every $n$-vertex graph containing no graph obtainable from $H$ by replacing each edge with a longer path of length $2\ell\pmod{k}$ and then possibly complementing as an induced subgraph, contains two sets $A,B$ of vertices such that $|A|,|B|\geqslant\epsilon n$ and $A$ is either complete or anticomplete to $B$.
For a vertex $v$ of a graph $G$, we let $N^{r}(v)$ denote the set of all vertices at distance exactly $r$ from $v$ in $G$. Similarly, we let $N^{r}[v]$ denote the set of all vertex at distance at most $r$ from $v$. When the graph $G$ is not clear from context, we use $N^{r}_{G}(v),N^{r}_{G}[v]$. For a set $X\subseteq V(G)$ of of vertices, we let $N^{r}[X]=\bigcup_{v\in X}N^{r}[v]$.
A massed graph $(G,\mu)$ is $(\delta,r)$-focused if for every $Z\subseteq V(G)$ with $\mu(Z)\geqslant\delta$, there is a vertex $v\in Z$ with $\mu(N^{r}_{G[Z]}[v])\geqslant\mu(Z)/2$.
For an integer $r\geqslant 1$, and $\epsilon>0$, we say that a massed graph $(G,\mu)$ is $(\epsilon,r)$-coherent if
  • $\mu(N^{r}[v])<\epsilon$ for every vertex $v$, and
  • $\min\{\mu(A),\mu(B)\}<\epsilon$ for every two anticomplete sets of vertices $A,B\subseteq V(G)$.
Note that every $(\epsilon,1)$-coherent graph is $\epsilon$-coherent.
The proof of Theorem 3.1 is split into the following two lemmas.
Lemma 3.7.
For every graph $H$, there exist $\epsilon>0$ and an integer $r\geqslant 1$, such that every $(\epsilon,r)$-coherent massed graph $(G,\mu)$ contains an induced proper fuzzy odd subdivision of $H$.
Lemma 3.8.
For every graph $H$ and integer $r\geqslant 1$, there exists $\epsilon,\delta>0$ such that every $(\delta,r)$-focused $\epsilon$-coherent massed graph $(G,\mu)$ contains an induced proper fuzzy odd subdivision of $H$.
Let us prove Theorem 3.1 assuming Lemma 3.7 and Lemma 3.8. This general strategy works for any hereditary class of graphs.
Proof of Theorem 3.1 assuming Lemma 3.7 and Lemma 3.8..
By Lemma 3.7 there exist $\epsilon_{r}>0$ and an integer $r\geqslant 1$, such that every $(\epsilon_{r},r)$-coherent massed graph $(G,\mu)$ contains an induced proper fuzzy odd subdivision of $H$. Choose $\epsilon_{r+1},\delta>0$ satisfying the hypotheses of Lemma 3.8 for $r+1$. Notice that $\epsilon_{r}$ and $\epsilon_{r+1}$ both could have been chosen to be arbitrarily small. So, by possibly choosing one or both of $\epsilon_{r}$ and $\epsilon_{r+1}$ to be smaller, we may assume that $\epsilon_{r}\delta=\epsilon_{r+1}$ and $\epsilon_{r+1}\leqslant\delta/2$.
Let $(G,\mu)$ be an $\epsilon_{r+1}$-coherent massed graph. By Lemma 3.8, we may assume that $(G,\mu)$ is not $(\delta,r+1)$-focused, since otherwise $G$ would contain an induced proper fuzzy odd subdivision of $H$. So, there exists $Z\subseteq V(G)$ with $\mu(Z)\geqslant\delta$, such that $\mu(N^{r+1}_{G[Z]}[v])<\mu(Z)/2$ for each $v\in Z$. Let $\mu^{\prime}(X)=\mu(X)/\mu(Z)$ for all $X\subseteq Z$. Then $(G[Z],\mu^{\prime})$ is a massed graph.
Suppose for sake of contradiction that $(G[Z],\mu^{\prime})$ is not $(\epsilon_{r},r)$-coherent. Note that $\mu(X)\geqslant\delta\mu^{\prime}(X)$ for all $X\subseteq Z$. In particular, if $\mu^{\prime}(X)\geqslant\epsilon_{r}$, then $\mu(X)\geqslant\epsilon_{r}\delta=\epsilon_{r+1}$. So since $(G,\mu)$ is $\epsilon_{r+1}$-coherent, $G[Z]$ contains no anticomplete pair $A,B\subseteq Z$ with $\mu^{\prime}(A),\mu^{\prime}(B)\geqslant\epsilon_{r}$. Hence, there exists $v\in Z$ such that $\mu^{\prime}(N_{G[Z]}^{r}[v])\geqslant\epsilon_{r}$. Therefore $\mu(N_{G[Z]}^{r}[v])\geqslant\epsilon_{r}\delta=\epsilon_{r+1}$. Since $\mu(N_{G[Z]}^{r+1}[v])<\mu(Z)/2$, we have that $\mu(Z\backslash N_{G[Z]}^{r+1}[v])>\mu(Z)/2\geqslant\delta/2\geqslant\epsilon_{r+1}$. This contradicts the fact that $(G,\mu)$ is $\epsilon_{r+1}$-coherent since $N_{G[Z]}^{r}[v]$, and $Z\backslash N_{G[Z]}^{r+1}[v]$ are anticomplete. Hence, $(G[Z],\mu^{\prime})$ is $(\epsilon_{r},r)$-coherent, and the result follows from Lemma 3.7. ∎
So now proving Theorem 3.1 (and thus Corollary 3.2, Theorem 1.1, Corollary 1.2, and Theorem 1.3) just relies on proving Lemma 3.7 and Lemma 3.8. We will prove Lemma 3.7 in Section 4, and Lemma 3.8 in Section 5. For both of these proofs, we shall need a lemma of Chudnovsky, Scott, Seymour, and Spirkl [3]. The rest of this section is dedicated to introducing this lemma. We require some further definitions first.
Let $\mathcal{Y}$ be a set of pairwise disjoint subsets of $V(G)$, where $G$ is a graph. Let $N$ be a graph, and for each $v\in V(N)$ let $X_{v}\subseteq V(G)$. We say that the family $X_{v}$ $(v\in V(N))$ is $\mathcal{Y}$-spread if for each $v\in V(N)$ there exists $Y_{v}\in\mathcal{Y}$ such that the sets $Y_{v}$ $(v\in V(N))$ are all different, and $X_{v}\subseteq Y_{v}$ for each $v\in V(N)$.
For disjoint $X,Y\subseteq V(G)$, we say that $X$ covers $Y$ if every vertex in $Y$ has a neighbour in $X$. Let $(G,\mu)$ be a massed graph. We say that $X\subseteq V(G)$ is $\delta$-dominant if $\mu(N[X])\geqslant\delta$.
A leaf of a tree $T$ is a vertex of degree at most 1. We let $L(T)$ denote the set of leaves of a tree $T$. A caterpillar is a tree $T$ such that $T-L(T)$ is a path. A rooted caterpillar is a tree $T$ with a distinguished vertex $h$, called its head, such that $T-L(T)$ is a path with one end being $h$.
Let $N$ be a rooted caterpillar with head $h$. Let $(G,\mu)$ be a massed graph. A $\delta$-realization of $N$ in $G$ is an assignment of a subset $X_{v}\subseteq V(G)$ to each vertex $v\in V(N)$, satisfying the following conditions:
  • the sets $X_{v}$ $(v\in V(N))$ are pairwise disjoint,
  • for every edge $uv$ of $N$, if $v$ lies on the path of $N$ between $u$ and $h$, then $X_{u}$ covers $X_{v}$,
  • for all distinct $u,v\in V(N)$, if $u,v$ are nonadjacent in $N$, then $X_{u},X_{v}$ are anticomplete, and
  • $\mu(X_{h})\geqslant\delta$, and for each $u\in V(N)\backslash\{h\}$, $X_{u}$ is connected and $\delta$-dominant.
The following lemma is due to Chudnovsky, Scott, Seymour, and Spirkl [3].
Lemma 3.9 (Chudnovsky, Scott, Seymour, and Spirkl [3, 5.2]).
Let $N$ be a rooted caterpillar, let $\delta,\epsilon>0$, let $(G,\mu)$ be an $\epsilon$-coherent massed graph, and let $\mathcal{Y}$ be a set of disjoint subsets of $V(G)$ such that $|\mathcal{Y}|=2^{|V(N)|}$ and $\mu(Y)\geqslant 2^{2^{|V(N)|}}(\delta+\epsilon)$ for each $Y\in\mathcal{Y}$. Then there is a $\mathcal{Y}$-spread $\delta$-realization of $N$ in $G$.
If in addition $r\geqslant 0$ is an integer, $\epsilon\leqslant\delta/2$, and $(G,\mu)$ is $(\delta,r)$-focused, then there is a $\mathcal{Y}$-spread $\delta$-realization $(X_{v}:v\in V(N))$ of $N$ in $G$ such that $X_{v}$ has an $r$-centre for each $v\in V(N)$, except the head.
The case that $(G,\mu)$ is $(\delta,r)$-focused is called the focused case of Lemma 3.9, and the other case is called the unfocused case. The unfocused case will be used in Section 4 to prove Lemma 3.7, and the focused case will be used in Section 5 to prove Lemma 3.8.
4 When small balls have small mass
This section is dedicated to proving Lemma 3.7. We require another lemma of Chudnovsky, Scott, Seymour, and Spirkl [3], but first we must introduce ladders.
A $k$-ladder in a graph $G$ is a family of $3k$ subsets
$A_{1},\ldots,A_{k},B_{1},\ldots,B_{k},C_{1},\ldots,C_{k}$
of $V(G)$, pairwise disjoint and such that
  • for $1\leqslant i\leqslant k$, $A_{i}$ is connected and covers $B_{i}$, and $B_{i}$ covers $C_{i}$,
  • for $1\leqslant i\leqslant k$, $A_{i},C_{i}$ are anticomplete, and
  • for all distinct $i,j\in\{1,\ldots,k\}$, $A_{i}$ is anticomplete to $A_{j}\cup B_{j}\cup C_{j}$.
A $k$-ladder is half-cleaned, if $B_{i}$ is anticomplete to $C_{j}$ for every $1\leqslant i<j\leqslant k$. The union of a $k$-ladder is the triple $(A,B,C)=A\cup B\cup C$, where $A=\bigcup_{i=1}^{k}A_{i}$, $B=\bigcup_{i=1}^{k}B_{i}$, and $C=\bigcup_{i=1}^{k}C_{i}$.
Chudnovsky, Scott, Seymour, and Spirkl [3] proved the following.
Lemma 4.1 (Chudnovsky, Scott, Seymour, and Spirkl [3, 6.2]).
Let $\epsilon,\kappa>0$, and let $k\geqslant 0$ be an integer such that $(k-1)k(\kappa+\epsilon)\leqslant 1$ and $(k-1)(\kappa+\epsilon)+4\epsilon\leqslant 1$. Let $(G,\mu)$ be an $\epsilon$-coherent massed graph. Then there is a half-cleaned $k$-ladder
$A_{1},\ldots,A_{k},B_{1},\ldots,B_{k},C_{1},\ldots,C_{k}$
in $G$ such that $\mu(C_{i})\geqslant\kappa$ for $1\leqslant i\leqslant k$.
The following lemma is implicit in the proof of [3, 6.4] due to Chudnovsky, Scott, Seymour, and Spirkl, and we closely follow their proof. This lemma is slightly stronger than we need, but the stronger statement better facilitates the proof.
Lemma 4.2.
Let $\epsilon,\kappa>0$, and let $N$ be a caterpillar with $k$ leaves $L$, with the property that ${(2^{|V(N)|}-1)2^{|V(N)|}(2^{2^{|V(N)|}}(\kappa+2\epsilon)+\epsilon)\leqslant 1}$ and $(2^{|V(N)|}-1)(2^{2^{|V(N)|}}(\kappa+2\epsilon)+\epsilon)+4\epsilon\leqslant 1$. Let $(G,\mu)$ be an $(\epsilon,5\cdot 2^{|V(N)|})$-coherent massed graph. Then there is an induced subdivided caterpillar $N^{\prime}$, which is isomorphic to a subdivision of $N$ with leaves $L^{\prime}=\{b_{1},\ldots,b_{k}\}$ and a half-cleaned $k$-ladder
$A_{1},\ldots,A_{k},B_{1},\ldots,B_{k},C_{1},\ldots,C_{k}$
in $G$, such that
  • $\mu(C_{i})\geqslant\kappa$ for $1\leqslant i\leqslant k$,
  • $N^{\prime}$ and $(A,B,C)$ are disjoint,
  • $V(N^{\prime})\backslash L^{\prime}$ and $(A,B,C)$ are anticomplete,
  • for each $i\in\{1,\ldots,k\}$, $N(b_{i})\cap(A,B,C)\subseteq A_{i}$, and
  • for each $i\in\{1,\ldots,k\}$, $G[A_{i}\cup\{b_{i}\}]$ is connected.
Proof.
Assign a head to a non-leaf vertex of $N$ to make it rooted. Let $k^{\prime}=2^{|V(N)|}$. By Lemma 4.1, there is a half-cleaned $k^{\prime}$-ladder
$A_{1},\ldots,A_{k^{\prime}},B_{1},\ldots,B_{k^{\prime}},C_{1},\ldots,C_{k^{\prime}}$
in $G$ such that $\mu(C_{i})\geqslant 2^{k^{\prime}}(\kappa+2\epsilon)$ for $1\leqslant i\leqslant k$. Then the unfocused case of Lemma 3.9 (taking $\delta=\kappa+\epsilon)$ implies that there is a $\{C_{1},\ldots,C_{k^{\prime}}\}$-spread $(\kappa+\epsilon)$-realization $(X_{v}:v\in V(N))$ of $N$ in $G$.
Let $n_{1},\ldots,n_{q}$ be the vertices of $N$ that are not leaves, where $n_{1}$ is the head of $N$ and $n_{i}n_{i+1}$ is an edge of $N$ for $1\leqslant i<q$. Choose $x_{1}\in X_{n_{1}}$. For $2\leqslant i\leqslant q$ in order, since $X_{n_{i}}$ covers $X_{n_{i-1}}$ by definition of a realization, we may choose a vertex $x_{i}\in X_{n_{i}}$ adjacent to $x_{i-1}$. Since $\{x_{1},\ldots,x_{q}\}\subseteq C_{1}\cup\cdots\cup C_{k^{\prime}}$, it follows that $\{x_{1},\ldots,x_{q}\}$ is anticomplete to $A_{1}\cup\cdots\cup A_{k^{\prime}}$. Also, since there are no edges between $X_{u},X_{v}$ for nonadjacent $u,v\in V(N)$, it follows that $x_{1},\ldots,x_{q}$ are the vertices in order of an induced path of $G$. Since $(X_{v}:v\in V(N))$ is a realization of $N$ in $G$, we have that for each leaf $\ell$ of $N$ with neighbour $n_{j}$ say, $x_{j}$ has a neighbour in $X_{\ell}$, and $X_{\ell}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}\backslash\{x_{j}\}$. Furthermore, for all distinct leaves $\ell,\ell^{\prime}$ of $N$, $X_{\ell}$ is anticomplete to $X_{\ell^{\prime}}$.
Let $I$ be the set of $i\in\{1,\ldots,k^{\prime}\}$ such that $X_{\ell}\subseteq C_{i}$ for some leaf $\ell$ of $N$. For each $i\in I$, let $\ell_{i}$ be the leaf of $N$ with $X_{\ell_{i}}\subseteq C_{i}$ (note that such a leaf is unique since $(X_{v}:v\in V(N))$ is $\{C_{1},\ldots,C_{k^{\prime}}\}$-spread. Let $i\in I$, and let $j\in\{1,\ldots,q\}$ such that $\ell_{i}$ is adjacent to $n_{j}$ in $N$. We define $x^{i}=x_{j}$. Thus there may be distinct values $i,i^{\prime}\in I$ with $x^{i}=x^{i^{\prime}}$.
Let $i\in I$. Since $X_{\ell_{i}}$ is $(\kappa+\epsilon)$-dominant, and $(G,\mu)$ is $(\epsilon,5k^{\prime})$-coherent, $N_{G}[X_{\ell_{i}}]\backslash N_{G}^{5k^{\prime}}[x_{1}]$ is non-empty. Therefore, there exists $u_{i}\in X_{\ell_{i}}$ with a neighbour $v$ such that the distance between $v$ and $x_{1}$ in $G$ is at least $15k^{\prime}+$ 1. Consequently, the distance between $u_{i}$ and $x_{1}$ in $G$ is at least $5k^{\prime}$. Since $X_{\ell_{i}}$ covers $X_{x^{i}}$, the induced subgraph $G[X_{\ell_{i}}\cup\{x^{i}\}]$ is connected. So there is a minimal path $P_{i}$ in $G[X_{\ell_{i}}\cup\{x_{i}\}]$ between $x^{i}$ and some vertex $u_{i}$ such that distance between $x_{1}$ and $u_{i}$ in $G$ is equal to $q+4i\leqslant 5k^{\prime}$. Choose a vertex $b_{i}\in B_{i}$ adjacent to $u_{i}$. Note that the distance between $x_{1}$ and $b_{i}$ in $G$ is between $q+4i-1$ and $q+4i+1$, so $\{b_{i}:i\in I\}$ is an independent set. Let $c_{i}$ be the second vertex of $P_{i}$, that is, the vertex adjacent to $x^{i}$, and let $Q_{i}$ be a shortest path of $G[V(P_{i})\cup\{b_{i}\}]\backslash\{x^{i}\}$ between $c_{i}$ and $b_{i}$. Thus $Q_{i}\backslash\{b_{i}\}$ is a path of $G[X_{\ell_{i}}]$. The subgraph of $G$ induced on $\{x_{1},\ldots,x_{q}\}\cup\{c_{i}:i\in I\}$ is a caterpillar isomorphic to $N$.
Next we will show that the subgraph $N^{\prime}$ of $G$ induced on $\{x_{1},\ldots,x_{q}\}\cup\bigcup_{i\in I}V(Q_{i})$ is a subdivided caterpillar isomorphic to a subdivision of $N$. Let $i\in I$. It is enough to show that $V(Q_{i})\backslash\{c_{i}\}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}$, and for every $j\in I$ with $i<j$, $V(Q_{i})$ is anticomplete to $V(Q_{j})$. Since $X_{\ell_{i}}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}\backslash\{x^{i}\}$, and $V(Q_{i})\backslash\{b_{i}\}\subseteq X_{\ell_{i}}$, we have that $V(Q_{i})\backslash\{b_{i}\}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}\backslash\{x^{i}\}$. By the minimality of $P_{i}$, $x^{i}$ has no neighbour in $V(Q_{i})\backslash\{b_{i}\}$. Furthermore, $b_{i}$ has no neighbour in $\{x_{1},\ldots,x_{q}\}$, since the distance in $G$ between $x_{1}$ and $b_{i}$ is at least $q+4i-1>q$, and $G[\{x_{1},\ldots,x_{q}\}]$ is a path. So, $V(Q_{i})\backslash\{c_{i}\}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}$. If $i$ is not the maximum element of $I$, then let $j\in I$ be such that $i<j$. Certainly $V(Q_{i})\backslash\{b_{i}\}$ and $V(Q_{j})\backslash\{b_{j}\}$ are anticomplete, since $V(Q_{i})\backslash\{b_{i}\}\subseteq X_{\ell_{i}}$, $V(Q_{j})\backslash\{b_{j}\}\subseteq X_{\ell_{j}}$, and $X_{\ell_{i}}$ and $X_{\ell_{j}}$ are anticomplete. Also $b_{i}$ has no neighbour in $V(Q_{j})\backslash\{b_{j}\}$ since the $k$-ladder is half-cleaned; so it remains to check that $b_{j}$ has no neighbour in $V(Q_{i})$. By the minimality of $P_{i}$, the distance in $G$ between $x_{1}$ and each vertex of $V(Q_{i})\backslash\{b_{i}\}$ is at most $q+4i$. So, the distance in $G$ between $x_{1}$ and each vertex of $V(Q_{i})$ is at most $q+4i+1$. Since, the distance in $G$ between $x_{1}$ and $b_{j}$ is at least $q+4j-1>q+4i+2$, it follows that $b_{j}$ has no neighbour in $V(Q_{i})$. Thus, $V(Q_{i})$ and $V(Q_{j})$ are anticomplete, and it follows that the subgraph $N^{\prime}$ of $G$ induced on $\{x_{1},\ldots,x_{q}\}\cup\bigcup_{i\in I}V(Q_{i})$ is a subdivided caterpillar isomorphic to a subdivision of $N$.
For $1\leqslant i\leqslant k^{\prime}$, let $C^{\prime}_{i}=C_{i}\backslash N_{G}^{q+4k^{\prime}+4}[x_{1}]$. Since $k^{\prime}\geqslant 2^{q+2}$, it follows that $45k^{\prime}\geqslant q+4k^{\prime}+$ 4, and so $(G,\mu)$ is $(\epsilon,q+4k^{\prime}+4)$-coherent. Hence $\mu(N_{G}^{q+4k^{\prime}+4}[x_{1}])<\epsilon$. Thus for $1\leqslant i\leqslant k^{\prime}$, $\mu(C^{\prime}_{i})\geqslant\mu(C_{i})-\epsilon\geqslant\kappa$. Since every vertex in $V(N^{\prime})$ has distance at most $q+4k^{\prime}+1$ from $x_{1}$ in $G$, it follows that every vertex in $C^{\prime}_{i}$ has distance at least three from $V(N^{\prime})$ in $G$. Let $B^{\prime}_{i}$ be the set of vertices in $B_{i}\backslash\{b_{i}\}$ with no neighbours in $V(N^{\prime})$. Since every vertex in $C^{\prime}_{i}$ has a neighbour in $B_{i}$ and has distance at least three from $V(N^{\prime})$ in $G$, it follows that $B_{i}^{\prime}$ covers $C^{\prime}_{i}$.
The sets
$A_{i}\ (i\in I),B_{i}^{\prime}\ (i\in I),C_{i}^{\prime}\ (i\in I)$
now form the desired $k$-ladder. ∎
For $I\subseteq[k]$, a $I$-tick of a $k$-ladder $(A,B,C)$ is a collection of induced paths $(Q_{i}:i\in I)$ of $G$, each with a distinct end $q_{i}$, and a common end $x_{I}$ such that
  • the sets $(Q_{i}\backslash\{x_{I}\}:i\in I)$ are pairwise anticomplete,
  • for each $i\in I$, $Q_{i}\backslash\{q_{i}\}$ is anticomplete to the $k$-ladder $(A,B,C)$,
  • for each $i\in I$, $\{q_{i}\}$ is disjoint to the $k$-ladder $(A,B,C)$, and $N(q_{i})\cap(A,B,C)\subseteq A_{i}$, and
  • for each $i\in I$, $A_{i}\cup\{q_{i}\}$ is connected.
We call $x_{I}$ the centre of the tick. Given a tick $(Q_{i}:i\in I)$ of a $k$-ladder $(A,B,C)$ with centre $x_{I}$, for each $i\in I$ and $c\in C_{i}$, let $P_{c}$ be a path in $G[V(Q_{i})\cup A_{i}\cup B_{i}\cup C_{i}]$ between $c$ and $x_{I}$ which is minimal subject to $V(P_{c})\cap C_{i}=\{c\}$ and $|V(P_{c})\cap B_{i}|=1$. If for every $i\in I$ and $c\in C_{i}$, the path $P_{c}$ has odd length, then we say that $(Q_{i}:i\in I)$ is an odd tick of the $k$-ladder $(A,B,C)$. Similarly, if for every $i\in I$ and $c\in C_{i}$, the path $P_{c}$ has even length, then we say that $(Q_{i}:i\in I)$ is an even tick of the $k$-ladder $(A,B,C)$.
Now, we derive the following from Lemma 4.2.
Lemma 4.3.
Let $\epsilon,\kappa>0$, such that $(2^{4pq+4q-1}-1)2^{4pq+4q-1}(2^{2^{4pq+4q-1}}(2\kappa+2\epsilon)+\epsilon)\leqslant 1$ and $(2^{4pq+4q-1}-1)(2^{2^{4pq+4q-1}}(2\kappa+2\epsilon)+\epsilon)+4\epsilon\leqslant 1$. Let $(G,\mu)$ be an $(\epsilon,5\cdot 2^{4pq+4q-1})$-coherent massed graph. Then there is a half-cleaned $pq$-ladder
$A_{1},\ldots,A_{pq},B_{1},\ldots,B_{pq},C_{1},\ldots,C_{pq}$
in $G$, with pairwise anticomplete ticks $(Q_{i}^{1}:i\in I_{1}),\ldots,(Q_{i}^{q}:i\in I_{q})$, such that
  • $\mu(C_{i})\geqslant\kappa$ for $1\leqslant i\leqslant pq$,
  • $I_{1}\cup\cdots\cup I_{q}=\{1,\ldots,pq\}$,
  • $|I_{j}|=p$ for each $1\leqslant j\leqslant q$, and
  • either every tick is even, or every tick is odd.
Proof.
First observe that a forest consisting of $2q$ stars, each with $2p$ leaves is an induced subgraph of a caterpillar with $14pq+4q-$ 1 vertices, and $4pq$ leaves. So, by Lemma 4.2, $(G,\mu)$ contains a half-cleaned $4pq$-ladder
$A_{1},\ldots,A_{4pq},B_{1},\ldots,B_{4pq},C_{1},\ldots,C_{4pq},$
with pairwise anticomplete ticks $(Q_{i}^{1}:i\in I_{1}),\ldots,(Q_{i}^{2q}:i\in I_{2q})$, such that
  • $\mu(C_{i})\geqslant 2\kappa$ for $1\leqslant i\leqslant 4pq$,
  • $I_{1}\cup\cdots\cup I_{2q}=\{1,\ldots,4pq\}$, and
  • $|I_{j}|=2p$ for each $1\leqslant j\leqslant 2q$.
Let $x_{1},\ldots,x_{2q}$ be the centres of the ticks $(Q_{i}^{1}:i\in I_{1}),\ldots,(Q_{i}^{2q}:i\in I_{2q})$. For each $j\in[2q]$, $i\in I_{j}$, and $c\in C_{i}$, let $P_{c}$ be a path in $G[V(Q_{i})\cup A_{i}\cup B_{i}\cup C_{i}]$ between $c$ and $x_{j}$ which is minimal subject to $V(P_{c})\cap C_{i}=\{c\}$ and $|V(P_{c})\cap B_{i}|=1$. For every $i\in[4pq]$, since $\mu(C_{i})\geqslant 2\kappa$, there exists $C_{i}^{\prime}\subseteq C_{i}$ such that $\mu(C_{i}^{\prime})\geqslant\mu(C_{i})/2\geqslant\kappa$, and either for every $c\in C^{\prime}_{i}$, the path $P_{c}$ has odd length, or for every $c\in C^{\prime}_{i}$, the path $P_{c}$ has even length. Let $C^{\prime}=\bigcup_{i\in[4pq]}C_{i}^{\prime}$.
Similarly, for every $j\in[2q]$, there exists a $I_{j}^{\prime}\subseteq I_{j}$ with $|I_{j}^{\prime}|=q$ such that either for every $i\in I_{j}^{\prime}$, and $c\in C_{i}^{\prime}$, the path $P_{c}$ has odd length, or for every $i\in I_{j}^{\prime}$, and $c\in C_{i}^{\prime}$, the path $P_{c}$ has even length. In other words, for every $j\in[2q]$, there exists a $I_{j}^{\prime}\subseteq I_{j}$ with $|I_{j}^{\prime}|=q$ such that the tick $(Q_{i}^{j}:i\in I_{j}^{\prime})$ is either odd or even.
Going one step further, there exists a $Q\subseteq[2q]$ with $|Q|=q$ such that either for every $j\in Q$, $(Q_{i}^{j}:i\in I_{j}^{\prime})$ is a odd tick of $(A,B,C^{\prime})$, or for every $j\in Q$, $(Q_{i}^{j}:i\in I_{j}^{\prime})$ is a even tick of $(A,B,C^{\prime})$. This gives us the desired ticks and $pq$-ladder. ∎
For a set $V$ and a non-negative integer $k$, we let (${V\choose k}$ denote the set of all $k$-element subsets of $k$. We are now ready to prove Lemma 3.7.
Proof of Lemma 3.7..
Clearly it is enough to prove the lemma in the case that $H$ is a complete graph. So, let $H=K_{n}$. Let $\epsilon=2^{-2^{4n^{2}}-4n^{2}}n^{-2}$, $r=5\cdot 2^{4n^{2}}$, and let $(G,\mu)$ be a $(\epsilon,r)$-coherent massed graph.
By Lemma 4.3, there is a half-cleaned $n(n-1)$-ladder
$A_{1},\ldots,A_{n(n-1)},B_{1},\ldots,B_{n(n-1)},C_{1},\ldots,C_{n(n-1)}$
in $G$, with pairwise anticomplete ticks $(Q_{i}^{1}:i\in I_{1}),\ldots,(Q_{i}^{n}:i\in I_{n})$, such that
  • $\mu(C_{i})\geqslant n(n-1)\epsilon$ for $1\leqslant i\leqslant n(n-1)$,
  • $I_{1}\cup\cdots\cup I_{n}=\{1,\ldots,n(n-1)\}$,
  • $|I_{j}|=n-1$ for each $1\leqslant j\leqslant n$, and
  • either every tick is even, or every tick is odd.
Let $x_{1},\ldots,x_{n}$ be the centres of the ticks $(Q_{i}^{1}:i\in I_{1}),\ldots,(Q_{i}^{n}:i\in I_{n})$. For each $j\in[n]$, $i\in I_{j}$, and $c\in C_{i}$, let $P_{c}$ be a path in $G[V(Q_{i})\cup A_{i}\cup B_{i}\cup C_{i}]$ between $c$ and $x_{j}$ which is minimal subject to $V(P_{c})\cap C_{i}=\{c\}$ and $|V(P_{c})\cap B_{i}|=1$. Note that either all such paths $P_{c}$ have even length, or they all have odd length. For each such $j\in[n]$, $i\in I_{j}$, and $c\in C_{i}$, let $b_{c}$ be the unique vertex of $V(P_{c})\cap B_{i}$. Observe that for each $j,j^{\prime}\in[n]$ with $j\not=j^{\prime}$, $i\in I_{j}$, $i^{\prime}\in I_{j^{\prime}}$ with $i<i^{\prime}$, $c\in C_{i}$ and $c^{\prime}\in C_{i^{\prime}}$ with $c$ adjacent to $c^{\prime}$, the subgraph $G[V(P_{c})\cup V(P_{c^{\prime}})]\backslash\{b_{c}b_{c^{\prime}},cb_{c^{\prime}}\}$ is an odd path between $x_{j}$ and $x_{j^{\prime}}$, and so in particular, one of $G[V(P_{c})\cup V(P_{c^{\prime}})]$ or $G[(V(P_{c})\cup V(P_{c^{\prime}}))\backslash\{c,c^{\prime}\}]$ is a fuzzy odd path between $x_{j}$ and $x_{j^{\prime}}$.
Arbitrarily order the elements (${[n]\choose 2}$ by $p_{1},\ldots,p_{\frac{1}{2}n(n-1)}$. For $1\leqslant m\leqslant\frac{1}{2}n(n-1)$, let $s_{m}\in I_{i}$, $s_{m}^{\prime}\in I_{j}$ be such that $\bigcup_{\{i,j\}\in{[n]\choose 2}}\{s_{m},s_{m}^{\prime}\}=[n(n-1)]$. For $1\leqslant m\leqslant\frac{1}{2}n(n-1)$, we will inductively find edges $c_{1}c_{1}^{\prime},\ldots,c_{m}c_{m}^{\prime}$ such that
  • for every $1\leqslant i\leqslant m$, $c_{i}\in C_{s_{i}}$, and $c_{i}^{\prime}\in C_{s_{i}^{\prime}}$, and
  • for every pair $1\leqslant i<j\leqslant m$, the distance in $G$ between $\{c_{i},c_{i}^{\prime}\}$ and $\{c_{j},c_{j}^{\prime}\}$ is at least 4.
Let $1\leqslant m\leqslant\frac{1}{2}n(n-1)$, and suppose we have already found such edges $c_{1}c_{1}^{\prime},\ldots,c_{m-1}c_{{m-1}}^{\prime}$. Since $(G,\mu)$ is $(\epsilon,3)$-coherent, for $s\in\{s_{m},s_{m}^{\prime}\}$ we have that
$\mu\left(C_{s}\backslash\bigcup_{i=1}^{m-1}N_{G}^{3}[\{c_{i},c_{{i}}^{\prime}\}]\right)\geqslant\mu(C_{s})-\mu\left(\bigcup_{i=1}^{m-1}N_{G}^{3}[\{c_{{i}},c_{{i}}^{\prime}\}]\right)>n(n-1)\epsilon-2(m-1)\epsilon>\epsilon.$
Therefore, since $(G,\mu)$ is $\epsilon$-coherent, there exists vertices $c_{m}\in C_{s_{m}}\backslash\bigcup_{i=1}^{m-1}N_{G}^{3}[\{c_{{i}},c_{{i}}^{\prime}\}]$, and $c_{m}^{\prime}\in C_{s_{m}^{\prime}}\backslash\bigcup_{i=1}^{m-1}N_{G}^{3}[\{c_{{i}},c_{{i}}^{\prime}\}]$, with $c_{m}$ and $c_{m}^{\prime}$ adjacent. Clearly the distance in $G$ between $\{c_{m},c_{m}^{\prime}\}$ and $\bigcup_{i=1}^{m-1}\{c_{{i}}c_{{i}}^{\prime}\}$ is at least 4. Hence $c_{1}c_{1}^{\prime},\ldots,c_{m}c_{m}^{\prime}$ provides the desired collection of edges. So, by induction, we can find such edges $c_{1}c_{1}^{\prime},\ldots,c_{{\frac{1}{2}n(n-1)}}c_{{\frac{1}{2}n(n-1)}}^{\prime}$.
For each $p_{a}=\{i,j\}\in{[n]\choose 2}$, observe that one of $G[V(P_{c_{a}})\cup V(P_{c_{a}^{\prime}})]$ or $G[(V(P_{c_{a}})\cup V(P_{c_{a}^{\prime}}))\backslash\{c_{a},c_{a}^{\prime}\}]$ is a fuzzy odd path between $x_{i}$ and $x_{j}$. So, for each $p_{a}=\{i,j\}\in{[n]\choose 2}$, choose $Q_{\{i,j\}}$ from $G[V(P_{c_{a}})\cup V(P_{c_{a}^{\prime}})]$ and $G[(V(P_{c_{a}})\cup V(P_{c_{a}^{\prime}}))\backslash\{c_{a},c_{a}^{\prime}\}]$ so that $Q_{\{i,j\}}$ is a fuzzy odd path between $x_{i}$ and $x_{j}$. Furthermore, for $\{i,j\},\{i^{\prime},j^{\prime}\}\in{[n]\choose 2}$ with $p_{a}=\{i,j\}$, $p_{a^{\prime}}=\{i^{\prime},j^{\prime}\}$, and $\{i,j\}\not=\{i^{\prime},j^{\prime}\}$, since the distance in $G$ between $\{c_{{a}},c_{a}^{\prime}\}$ and $\{c_{a^{\prime}},c_{a^{\prime}}^{\prime}\}$ is at least 4, we have that $V(Q_{\{i,j\}})\backslash\{x_{i},x_{j}\}$ is anticomplete to $V(Q_{\{i^{\prime},j^{\prime}\}})\backslash\{x_{i^{\prime}},x_{j^{\prime}}\}$. Therefore, $G[\bigcup_{\{i,j\}\in{[n]\choose 2}}V(Q_{\{i,j\}})]$ is a proper fuzzy odd subdivision of $K_{n}$ as desired. ∎
5 Focused massed graphs
This section is dedicated to proving Lemma 3.8.
Let $T^{\prime}$ be a caterpillar and let $(G,\mu)$ be a massed graph containing an induced subgraph $T$ isomorphic to $T^{\prime}$. Let $x_{1},\ldots,x_{q}$ be the vertices of $T$ with degree more than one, and for each leaf $v$ of $T$, let $x^{v}$ be its neighbour in $\{x_{1},\ldots,x_{q}\}$. Let $\kappa>0$, and let $r\geqslant 1$ be an integer. If for each leaf $v$ of $T$ there is a subset $X_{v}\subseteq V(G)$ with the follow properties;
  • for each leaf $v\in L(T)$, $v\in X_{v}$, $x^{v}\not\in X_{v}$, and $X_{v}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}\backslash\{x^{v}\}$,
  • for all distinct leaves $u,v\in L(T)$, $X_{u}$ is anticomplete to $X_{v}$,
  • for each leaf $v\in L(T)$, $x^{v}$ is an $r$-centre for $X_{v}\cup\{x^{v}\}$,
  • for each leaf $v\in L(T)$, $v$ is the unique neighbour of $x^{v}$ in $X_{v}$, and
  • for each leaf $v\in L(T)$, $X_{v}\cup\{x^{v}\}$ is $\kappa$-dominant,
then we say that $(T,(X_{v})_{v\in L(T)})$ is a $(T^{\prime},r,\kappa)$-frame of $(G,\mu)$.
A result very similar to the following lemma is implicit in the proof of [3, 7.2] due to Chudnovsky, Scott, Seymour, and Spirkl. Since it is not given explicitly, we give a proof closely following [3, 7.2].
Lemma 5.1.
Let $T^{\prime}$ be a caterpillar with $t\geqslant 3$ vertices, and let $r\geqslant 1$ be an integer. Let $\epsilon,\delta,\kappa>0$ with $t\epsilon\leqslant\delta\leqslant\kappa\leqslant 2^{-2^{t}}t^{-2^{t}}(2^{-2^{t}}(2^{-t}-\epsilon)-\epsilon)$. Then, every $(\delta,r)$-focused $\epsilon$-coherent massed graph $(G,\mu)$ contains a $(T^{\prime},3r+2,\kappa)$-frame.
Proof.
We may assume that $T^{\prime}$ is rooted, and its head is an internal vertex. Let $k=2^{t}$, and let $\lambda=1/k-\epsilon$.
For each $1\leqslant j\leqslant k$ in order, choose some $Y_{j}\subset V(G)\backslash\left(\bigcup_{i<j}Y_{i}\right)$ minimal, subject to $\mu(Y_{j})\geqslant\lambda$ (so with $\mu(Y_{j})<\lambda+\epsilon)$. This is possible since
$\mu\left(V(G)\backslash\left(\bigcup_{i<j}Y_{i}\right)\right)$ $\geqslant 1-\mu\left(\bigcup_{i<j}Y_{i}\right)$
$\geqslant 1-\sum_{i<j}\mu(Y_{i})$
$>1-(j-1)(\lambda+\epsilon)$
$=1-(j-1)/k$
$\geqslant 1/k$
$>\lambda.$
So, $Y_{1},\ldots,Y_{k}$ are disjoint, and $\mu(Y_{j})\geqslant\lambda$ for all $1\leqslant j\leqslant k$.
For $0\leqslant i\leqslant k$, let $\kappa_{i}=2^{-i}t^{-i}(2^{-k}\lambda-\epsilon)$. Then, $\kappa_{0}\geqslant\kappa_{k}=2^{-2^{t}}t^{-2^{t}}(2^{-2^{t}}(2^{-t}-\epsilon)-\epsilon)\geqslant\kappa\geqslant\delta$. As $\kappa_{0}\geqslant\delta$ and $G$ is $(\delta,r)$-focused, $G$ is also $(\kappa_{0},r)$-focused. So by Lemma 3.9, there is a $\{Y_{1},\ldots,Y_{k}\}$-spread $\kappa_{0}$-realization $(X_{v}:v\in V(T^{\prime}))$ of $T^{\prime}$ in $G$ such that $X_{v}$ has an $r$-centre for each $v\in V(T^{\prime})$, except the head. Let the vertices of degree more than one in $T^{\prime}$ be $t_{1},\ldots,t_{q}$, where $t_{1}$ is the head, and $t_{i}t_{i+1}$ is an edge for each $11\leqslant i\leqslant q-$ 1. Choose $x_{1}\in X_{t_{1}}$, and inductively for each $2\leqslant i\leqslant q$, choose $x_{i}\in X_{t_{i}}$ such that $x_{i}$ is adjacent to $x_{i-1}$. Then, $x_{1},\ldots,x_{q}$ are the vertices of an induced path in $G$.
Let $L(T^{\prime})=\{v_{1},\ldots,v_{\ell}\}$, and for each $v\in L(T^{\prime})$, let $x^{v}$ be the vertex $x_{j}\in\{x_{1},\ldots,x_{q}\}$ such that $v$ is adjacent to $t_{j}$ in $T^{\prime}$. Since $X_{v}$ has an $r$-centre, and $x^{v}$ has a neighbour in $X_{v}$, it follows that $x^{v}$ is a $(2r+1)$-centre of $X_{v}\cup\{x^{v}\}$.
For each $v\in L(T^{\prime})$, let $X^{0}_{v}=X_{v}\cup\{x^{v}\}$. For each $0\leqslant i\leqslant\ell$ in order, we shall inductively define $X^{i}_{v}$ for each $v\in L(T^{\prime})$ satisfying the following:
  • for each $v\in L(T^{\prime})$, $x^{v}\in X_{v}^{i}$ and $X_{v}^{i}\backslash\{x^{v}\}$ is anticomplete to $\{x_{1},\ldots,x_{q}\}\backslash\{x^{v}\}$,
  • for all distinct $u,v\in L(T^{\prime})$, $X_{u}^{i}\backslash\{x^{u}\}$ is anticomeplte to $X_{v}^{i}\backslash\{x^{v}\}$,
  • for $1\leqslant j\leqslant i$, $x^{v_{j}}$ is a $(3r+2)$-centre of $X^{i}_{v_{j}}$, and $x^{v_{j}}$ has a unique neighbour in $X^{i}_{v_{j}}$,
  • for $i<j\leqslant\ell$, $x^{v_{j}}$ is a $(2r+1)$-centre of $X^{i}_{v_{j}}$, and
  • for each $v\in L(T^{\prime})$, $X^{i}_{v}$ is $\kappa_{i}$-dominant.
The sets $(X^{0}_{v}:v\in L(T^{\prime}))$ satisfy these conditions. So, suppose that $1\leqslant i\leqslant\ell$, and we have defined $X^{i-1}_{v}$ for each $v\in V(T^{\prime})$ as above.
Let $v_{j}\in L(T^{\prime})\backslash\{v_{i}\}=\{v_{1},\ldots,v_{\ell}\}\backslash\{v_{i}\}$. If $j<i$, then let $r_{j}=3r+2$ if $j<i$, otherwise if $j>i$, then let $r_{j}=2r+1$. So in either case $x^{v_{j}}$ is an $r_{j}$-centre of $X^{i-1}_{v_{j}}$. Choose $X^{i}_{v_{j}}\subseteq X^{i-1}_{v_{j}}$, minimal such that $X^{i}_{v_{j}}$ is $\kappa_{i}$-dominant and $x^{v_{j}}$ is an $r_{j}$-centre for $X^{i}_{v_{j}}$. It follows from the minimality that $\mu(N_{G}[X^{i}_{v_{j}}])\leqslant\kappa_{i}+\epsilon$.
Since $X^{i-1}_{v_{i}}$ is $\kappa_{i-1}$-dominant, we have that $\mu(N_{G}[X^{i-1}_{v_{i}}])\geqslant\kappa_{i-1}$. Let
$Y^{\prime}=N_{G}[X^{i-1}_{v_{i}}]\backslash\left(N_{G}[\{x_{1},\ldots,x_{q}\}]\cup\bigcup_{u\in L(T^{\prime})\backslash\{v_{i}\}}N_{G}[X^{i}_{u}]\right).$
Then,
$\mu(Y^{\prime})$ $\geqslant\mu(N_{G}[X^{i-1}_{v_{i}}])-\mu(N_{G}[\{x_{1},\ldots,x_{q}\}])-\left(\sum_{{u\in L(T^{\prime})\backslash\{v_{i}\}}}\mu(N_{G}[X^{i}_{u}])\right)$
$\geqslant\kappa_{i-1}-q\epsilon-(\ell-1)(\kappa_{i}+\epsilon)$
$=2t\kappa_{i}-q\epsilon-(\ell-1)(\kappa_{i}+\epsilon)$
$=(2t-\ell+1)\kappa_{i}-(q+\ell-1)\epsilon$
$\geqslant 3\kappa_{i}-t\epsilon$
$\geqslant 2\kappa_{i}$
$\geqslant 2\kappa$
$\geqslant\delta.$
Since $G$ is $(\delta,r)$-focused, there is a subset $Y\subseteq Y^{\prime}$ with $\mu(Y)\geqslant\mu(Y^{\prime})/2\geqslant\kappa_{i}$, such that $Y$ has an $r$-centre $y$ say. Since $x^{v_{i}}$ is a $(2r+1)$-centre for $X^{i-1}_{v_{i}}$, and $y\in N_{G}[X^{i-1}_{v_{i}}]$, there is an induced path $P$ of $G[X^{i-1}_{v}\cup\{y\}]$ between $x^{v_{i}}$ and $y$, of length at most $22r+$ 2. Let $X^{i}_{v_{i}}=Y\cup V(P)$. Then $x^{v_{i}}$ is a $(3r+2)$-centre for $X^{i}_{v_{i}}$. Since $x^{v_{i}}$ is anticomplete to $Y$ and has only one neighbour in $V(P)$, it follows that $x^{v_{i}}$ has only one neighbour in $X^{i}_{v_{i}}$. Moreover, $X^{i}_{v_{i}}$ is $\kappa_{i}$-dominant since $\mu(X^{i}_{v_{i}})\geqslant\mu(Y)\geqslant\kappa_{i}$. This completes the inductive definition.
For each $v\in L(T^{\prime})$, let $c_{v}$ be the unique neighbour of $x^{v}$ in $X^{\ell}_{v}$. The subgraph $T$ induced on $\{x_{1},\ldots,x_{q}\}\cup\{c_{v}:v\in L(T^{\prime})\}$ is isomorphic to $T^{\prime}$. For each $c_{v}\in L(T)$, let $X_{c_{v}}=X_{v}^{\ell}$. Then, $(T,(X_{c_{v}})_{c_{v}\in L(T)})$ is a $(T^{\prime},3r+2,\kappa)$-frame of $G$, as desired. ∎
Before proving Lemma 3.8, we require one final lemma that allows us to find certain induced fuzzy odd paths.
Lemma 5.2.
Let $r\geqslant 1$ be an integer, and let $\epsilon>0$. Let $(G,\mu)$ be an $\epsilon$-coherent massed graph with $u\in X_{u}\subseteq A\subseteq V(G)$ and $v\in X_{v}\subseteq B\subseteq V(G)$ such that
  • $u$ is an $r$-centre of $X_{u}$ and $v$ is an $r$-centre of $X_{v}$,
  • $X_{u}$ is anti-complete to $X_{v}$,
  • $A\subseteq N[X_{u}]$ and $B\subseteq N[X_{v}]$, and
  • $\mu(A),\mu(B)\geqslant(r(r+1)(2^{2r+3}+3)+1)\epsilon$.
Then $G$ contains an induced fuzzy odd path $P$ between $u$ and $v$ of length at most $32r+$ 3, such that $V(P)\subseteq A\cup B$.
Proof.
Since $(G,\mu)$ is $\epsilon$-coherent and $X_{u}$ is anti-complete to $X_{v}$, we may assume without loss of generality that $\mu(X_{u})<\epsilon$.
For $0\leqslant j\leqslant r$, let $L_{A}^{j}$ be the set of vertices in $X_{u}$ whose distance in $G[X_{u}]$ from $u$ is equal to $j$. For $0\leqslant j\leqslant r$, let $D_{A}^{j}=(A\backslash X_{u})\cap(N(L_{A}^{j})\backslash(\bigcup_{0\leqslant p<j}N(L_{A}^{p}))$. So, $\mu(D_{A}^{0})+\cdots+\mu(D_{A}^{r})\geqslant\mu(D_{A}^{0}\cup\cdots\cup D_{A}^{r})\geqslant\mu(A)-\mu(X_{u})>(r(r+1)(2^{2r+3}+3)+1)\epsilon-\epsilon=r(r+1)(2^{2r+3}+3)\epsilon$. Choose $j_{A}$ minimum such that $\mu(D_{A}^{j_{A}})\geqslant(2^{2r+3}+3)\epsilon$. Since $(G,\mu)$ is $\epsilon$-coherent, we have that $j_{A}\geqslant 1$. By the minimality of $j_{A}$, we have $\mu(\bigcup_{0\leqslant p<j_{A}}D^{p}_{A})\leqslant\mu(D_{A}^{0})+\cdots+\mu(D_{A}^{j_{A}})<j_{A}(2^{2r+3}+3)\epsilon\leqslant r(2^{2r+3}+3)\epsilon$.
Let $B^{\prime}=B\backslash(\bigcup_{0\leqslant p<j_{A}}D^{p}_{A})$, then
$\mu(B^{\prime})\geqslant\mu(B)-\mu\left(\bigcup_{0\leqslant p<j_{A}}D^{p}_{A}\right)>(r(r+1)(2^{2r+3}+3)+1)\epsilon-r(2^{2r+3}+3)\epsilon\geqslant(3r+4)\epsilon.$
For $0\leqslant j\leqslant r$, let $L_{B}^{j}$ be the set of vertices in $X_{v}$ whose distance in $G[X_{v}]$ from $v$ is equal to $j$. For $0\leqslant j\leqslant r$, let $D_{B}^{j}=(B^{\prime}\backslash X_{v})\cap(N(L_{B}^{j})\backslash(\bigcup_{0\leqslant p<j}N(L_{B}^{p}))$.
Suppose that there exists a $0\leqslant j_{B}\leqslant r$ such that $\mu(N(L_{B}^{j_{B}})\cap D_{A}^{j_{A}})\geqslant 4^{j_{B}}\epsilon$. Choose such a $j_{B}$ to be minimum, and let $Z=(N(L_{B}^{j_{B}})\cap D_{A}^{j_{A}})\backslash\left(\bigcup_{0\leqslant p<j_{B}}N(L_{B}^{p})\right)$. Since $(G,\mu)$ is $\epsilon$-coherent, we have that $j_{B}\geqslant 1$. Then,
$\mu(Z)$ $\geqslant\mu(N(L_{B}^{j_{B}})\cap D_{A}^{j_{A}})-\left(\sum_{p=0}^{j_{B}-1}\mu(N(L_{B}^{p})\cap D_{A}^{j_{A}})\right)$
$>4^{j_{B}}\epsilon-\sum_{p=0}^{j_{B}-1}4^{p}\epsilon$
$=4^{j_{B}}\epsilon-\frac{1}{3}(4^{j_{B}}-1)\epsilon$
$\geqslant 3\epsilon.$
Now, choose $L\subseteq L_{B}^{j_{B}}$ minimal such that $\mu(N(L)\cap Z)\geqslant\epsilon$. Then $\mu(N(L)\cap Z)<2\epsilon$, and therefore $\mu(Z\backslash N(L))\geqslant\mu(Z)-\mu(N(L)\cap Z)>3\epsilon-2\epsilon=\epsilon$. So, since $(G,\mu)$ is $\epsilon$-coherent, there exists adjacent vertices $b_{1},b_{2}$, with $b_{1}\in Z\backslash N(L)$ and $b_{2}\in N(L)\cap Z$. For $i\in\{1,2\}$, let $a_{i}$ be the vertex of $L_{A}^{j_{A}}$ adjacent to $b_{i}$, and let $P_{i}$ be a shortest path in $G[X_{u}]$ between $u$ and $a_{i}$. Then $P_{1}$ and $P_{2}$ both have length $j_{A}$. Let $c$ be a vertex of $L$ adjacent to $b_{2}$, and let $P_{c}$ be a shortest path in $G[X_{v}]$ between $v$ and $c$. Then, $P_{c}$ has length $j_{B}$. Let $P_{1}^{\prime}=G[V(P_{1})\cup\{b_{1},b_{2}\}\cup V(P_{c})]$, and $P_{2}^{\prime}=G[V(P_{2})\cup\{b_{2}\}\cup V(P_{c})]$. Clearly $V(P_{1}^{\prime}),V(P_{2}^{\prime})\subseteq A\cup B$. Also, $P_{1}^{\prime}\backslash\{a_{1}b_{2}\}$ is a path between $u$ and $v$ of length $j_{A}+j_{B}+3\leqslant 2r+3$ and $P_{2}^{\prime}$ is a path between $u$ and $v$ of length $j_{A}+j_{B}+2\leqslant 2r+2$. Therefore one of $P_{1}^{\prime}$ or $P_{2}^{\prime}$ provides the desired induced fuzzy odd path.
So, we may now assume that $\mu(N(L_{B}^{j})\cap D_{A}^{j_{A}})<4^{j}\epsilon$ for each $0\leqslant j\leqslant r$. Therefore, $\mu(N(X_{v})\cap D_{A}^{j_{A}})<\sum_{j=0}^{r}4^{j}\epsilon=\frac{1}{3}(4^{r+1}-1)\epsilon<2^{2r+3}\epsilon$. Let $Z=D_{A}^{j_{A}}\backslash N(X_{v})$. Then $\mu(Z)\geqslant\mu(D_{A}^{j_{A}})-\mu(N(X_{v})\cap D_{A}^{j_{A}})\geqslant(2^{2r+3}+3)\epsilon-2^{2r+3}\epsilon=3\epsilon$. Since $Z$ is anticomplete to $X_{v}$, we have that $\mu(X_{v})<\epsilon$. So, $\mu(D_{B}^{0})+\cdots+\mu(D_{B}^{r})\geqslant\mu(D_{B}^{0}\cup\cdots\cup D_{B}^{r})\geqslant\mu(B^{\prime})-\mu(X_{v})\geqslant(3r+4)\epsilon-\epsilon=3(r+1)\epsilon$. Now, choose $0\leqslant j_{B}\leqslant r$ minimum such that $\mu(D_{B}^{j_{B}})\geqslant 3\epsilon$. Choose $D\subseteq D_{B}^{j_{B}}$ minimal such that $\mu(N(D)\cap Z)\geqslant\epsilon$. Then $\mu(N(D)\cap Z)<2\epsilon$, and therefore $\mu(Z\backslash N(D))\geqslant\mu(Z)-\mu(N(D)\cap Z)>3\epsilon-2\epsilon=\epsilon$. So, since $(G,\mu)$ is $\epsilon$-coherent, there exists adjacent vertices $b_{1},b_{2}$, with $b_{1}\in Z\backslash N(D)$ and $b_{2}\in N(D)\cap Z$. For $i\in\{1,2\}$, let $a_{i}$ be the vertex of $L_{A}^{j_{A}}$ adjacent to $b_{i}$, and let $P_{i}$ be a shortest path in $G[X_{u}]$ between $u$ and $a_{i}$. Then $P_{1}$ and $P_{2}$ both have length $j_{A}$. Let $c_{1}$ be a vertex of $D$ adjacent to $b_{2}$, and let $c_{2}$ be a vertex of $L_{B}^{j_{B}}$ adjacent to $c_{2}$. Let $P_{c}$ be a shortest path in $G[X_{v}]$ between $v$ and $c_{2}$. Then, $P_{c}$ has length $j_{B}$. Let $P_{1}^{\prime}=G[V(P_{1})\cup\{b_{1},b_{2},c_{1}\}\cup V(P_{c})]$, and $P_{2}^{\prime}=G[V(P_{2})\cup\{b_{2},c_{1}\}\cup V(P_{c})]$. Clearly $V(P_{1}^{\prime}),V(P_{2}^{\prime})\subseteq A\cup B$. Also, $P_{1}^{\prime}\backslash\{a_{1}b_{2}\}$ is a path between $u$ and $v$ of length $j_{A}+j_{B}+4\leqslant 2r+4$ and $P_{2}^{\prime}$ is a path between $u$ and $v$ of length $j_{A}+j_{B}+3\leqslant 2r+3$. Therefore one of $P_{1}^{\prime}$ or $P_{2}^{\prime}$ provides the desired induced fuzzy odd path. ∎
We are now ready to prove Lemma 3.8. The rest of the proof adapts that of another lemma of Chudnovsky, Scott, Seymour, and Spirkl [3, 7.1].
Proof of Lemma 3.8..
Clearly it is enough to prove the lemma in the case that $H$ is a complete graph on $n\geqslant 3$ vertices. So, let $H=K_{n}$, and let $r\geqslant 1$ be an integer. Now let $\epsilon,\delta,\kappa>0$ be such that
  • $n^{2}\epsilon\leqslant\delta\leqslant\kappa\leqslant 2^{-2^{n^{2}}}n^{-2^{n^{2}}}(2^{-2^{n^{2}}}(2^{-n^{2}}-\epsilon)-\epsilon)$, and
  • $2^{\frac{1}{2}n(n-1)}n^{n(n-1)}\left(3n^{2}(r+3)\epsilon+((3r+2)(3r+3)(2^{6r+7}+3)+1)\epsilon\right)\leqslant\kappa$.
Such $\epsilon,\delta,\kappa>0$ can clearly be chosen when $\epsilon$ is sufficiently small. Then, let $(G,\mu)$ be a $(\delta,r)$-focused $\epsilon$-coherent massed graph.
For every $i\in[n]$, let $S_{i}^{\prime}$ be a star with centre $w_{i}^{\prime}$ and leaves $L(S_{i}^{\prime})=\{u_{\{i,j\}}^{\prime}:j\in[n]\backslash[i]\}\cup\{v_{\{j,i\}}^{\prime}:j\in[i-1]\}$. Let $T^{\prime}$ be a caterpillar with $n(n-1)+n-1=n^{2}-1$ vertices that contains the disjoint union of $S_{1}^{\prime},\ldots,S_{n}^{\prime}$ as an induced subgraph. Then by Lemma 5.1, $(G,\mu)$ contains a $(T^{\prime},3r+2,\kappa)$-frame $(T,(X_{v})_{v\in L(T)})$. For each $i\in[n]$, let $S_{i}$ be the induced star of $T$ corresponding to the induced star $S_{i}^{\prime}$ of $T^{\prime}$. Let $w_{i}$ be the centre of $S_{i}$ and let $L(S_{i})=\{u_{\{i,j\}}:j\in[n]\backslash[i]\}\cup\{v_{\{j,i\}}:j\in[i-1]\}$ for each $i\in[n]$. For each $i\in[n]$, and each leaf $w\in L(S_{i})$ of $S_{i}$, let $x^{w}=w_{i}$.
Let $\kappa_{0}=\kappa$. For each $1\leqslant i\leqslant\frac{1}{2}n(n-1)$ in order, let
$\kappa_{i}=\frac{1}{n^{2}}\left(\kappa_{i-1}-3n^{2}(r+3)\epsilon-((3r+2)(3r+3)(2^{6r+7}+3)+1)\epsilon\right).$
For each $1\leqslant i\leqslant\frac{1}{2}n(n-1)$, we have that $\kappa_{i}\geqslant\frac{1}{2n^{2}}\kappa_{i-1}$. So, $\kappa_{0}>\cdots>\kappa_{\frac{1}{2}n(n-1)}\geqslant 3n^{2}(r+3)\epsilon+((3r+2)(3r+3)(2^{6r+7}+3)+1)\epsilon$. Arbitrarily order the elements of (${[n]\choose 2}$ by $p_{1},\ldots,p_{\frac{1}{2}n(n-1)}$. For each $w\in L(T)$, let $X^{0}_{w}=X_{w}\cup\{x^{w}\}$. For $1\leqslant i\leqslant\frac{1}{2}n(n-1)$ in order, we define $X^{i}_{w}$ $(w\in L(T)\backslash\{u_{p_{1}},\ldots,u_{p_{i-1}},v_{p_{1}},\ldots,v_{p_{i-1}}\}))$ and $P_{i}$ inductively as follows. We assume $P_{1},\ldots,P_{i-1}$ and $X^{i-1}_{w}$ $(w\in L(T)\backslash\{u_{p_{1}},\ldots,u_{p_{i-1}},v_{p_{1}},\ldots,v_{p_{i-1}}\}))$ have been defined, such that
  • for $11\leqslant h\leqslant i-$ 1, $P_{h}$ is an induced fuzzy odd path between $u_{p_{h}}$ and $v_{p_{h}}$, of length at most $76r+$ 7,
  • for $11\leqslant h\leqslant i-$ 1, $V(P_{h})\backslash\{u_{p_{h}},v_{p_{h}}\}$ is anticomplete to $V(T)\backslash\{u_{p_{h}},v_{p_{h}}\}$,
  • for $11\leqslant h<h^{\prime}\leqslant i-$ 1, $V(P_{h})$ is anticomplete to $V(P_{h^{\prime}})$,
  • for $11\leqslant h\leqslant i-$ 1 and $v\in\{u_{p_{i}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$, $V(P_{h})$ is anticomplete to $X^{i-1}_{v}$,
  • for $w\in\{u_{p_{i}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$, $X^{i-1}_{w}$ is $\kappa_{i-1}$-dominant, and
  • for $w\in\{u_{p_{i}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$, $\{x^{w},w\}\subseteq X^{i-1}_{w}$ and $x^{w}$ is a $(3r+2)$-centre for $X^{i-1}_{w}$.
For each $w\in\{u_{p_{i+1}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i+1}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$, choose $X^{i}_{w}\subseteq X^{i-1}_{w}$, minimal such that $X^{i}_{w}$ is $\kappa_{i}$-dominant and $x^{w}$ is a $(3r+2)$-centre for $X^{i}_{w}$. By deleting a vertex in $X^{i}_{w}$ with maximum $G[X^{i}_{w}]$-distance from $x^{w}$, the minimality of $X^{i}_{w}$ implies that $\mu(N_{G}[X^{i}_{w}])\leqslant\kappa_{i}+\epsilon$. We also have that
$\mu\left(N_{G}\left[\{w_{1},\ldots,w_{n}\}\cup\bigcup_{h<i}V(P_{h})\right]\right)$ $\leqslant(n+(i-1)(6r+8))\epsilon$
$\leqslant(n+\frac{1}{2}n(n-1)(6r+8))\epsilon$
$<3n^{2}(r+2)\epsilon.$
Let $u=u_{p_{i}}$ and $v=v_{p_{i}}$. Then, let $Y$ be the set of all vertices at distance at most one from $\{w_{1},\ldots,w_{n}\}\cup\bigcup_{h<i}V(P_{h})$ and $X^{i}_{w}$ for each $w\in\{u_{p_{i+1}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i+1}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$. Observe that,
$\mu(Y)\leqslant 3n^{2}(r+2)\epsilon+n(n-1)(\kappa_{i}+\epsilon)<3n^{2}(r+3)\epsilon+n^{2}\kappa_{i}.$
Let $A=N_{G}[X^{i-1}_{u}]\backslash Y$, and $B=N_{G}[X^{i-1}_{v}]\backslash Y$. Then, since $X^{i-1}_{u}$ is $\kappa_{i-1}$-dominant, it follows that
$\mu(A)\geqslant\mu(N_{G}[X^{i-1}_{u}])-\mu(Y)>\kappa_{i-1}-3n^{2}(r+3)\epsilon-n^{2}\kappa_{i}\geqslant((3r+2)(3r+3)(2^{6r+7}+3)+1)\epsilon.$
Similarly, $\mu(B)>((3r+2)(3r+3)(2^{6r+7}+3)+1)\epsilon$. So, by Lemma 5.2 (with $r$ replaced by $23r+$ 2), there exists an induced fuzzy odd path $P_{i}$ between $u$ and $v$ of length at most $76r+$ 7 and with $V(P_{i})\subseteq A\cup B$. Since $V(P_{i})\subseteq A\cup B\subseteq N_{G}[X_{u}^{i-1}]\cup N_{G}[X_{v}^{i-1}]$, $P_{i}$ is anticomplete to $V(T)\backslash\{u,v\}$, $V(P_{h})$ for each $11\leqslant h\leqslant i-$ 1, and $X_{w}^{i}\subseteq X_{w}^{i-1}$ for each $w\in\{u_{p_{i}},\ldots,u_{p_{\frac{1}{2}n(n-1)}},v_{p_{i}},\ldots,v_{p_{\frac{1}{2}n(n-1)}}\}$. Therefore, $P_{i}$ provides the desired induced fuzzy odd path. This completes the inductive definition of $P_{1},\ldots,P_{\frac{1}{2}n(n-1)}$.
It now follows that $G[\{w_{1},\ldots w_{n}\}\cup\bigcup_{j=1}^{\frac{1}{2}n(n-1)}V(P_{j})]$ is a proper fuzzy odd subdivision of $H$, as desired. ∎
This completes the proof of Theorem 3.1, and thus also of Corollary 3.2, Theorem 1.1, Corollary 1.2, and Theorem 1.3.
Acknowledgements
The author thanks Jim Geelen and Rose McCarty for a number of helpful discussions on pivot-minors. The author also thanks the anonymous referees for helpful comments, in particular for spotting a serious error in an earlier version of this manuscript.